the length of the rectangle is 30 inches the width is 10 inches shorter than the land what is the area of the rectangle
Answer by amarjeeth123(569)

Length of the rectangle=30 inches
Width of the rectangle=30-10=20 inches
Area of the rectangle=Length*width=30(20)=600 square inches
Answer=600 square inches.